OpenStudy (anonymous):

In a group of 15 high school students 7 play football, 8 play baseball and 3 do neither. How many students play both football and baseball?

OpenStudy (anonymous):

15(all)-3(do neither)=12 play games 7+8-12=3 play both football and baseball

OpenStudy (anonymous):

15-3=12 now out of these 12 stdnts 7 play football and 8 play basketball. common is 3